cos a + cos b + cos c = sin a +sin b+ sin c =0 then value of cos 3a + cos 3b+ cos 3c is....Please show steps

Grade:

1 Answers

Deepak Kumar Shringi
5 years ago
cos3a+cos3b+cos3c
=4cos3a-3cosa+4cos3b-3cosb+4cos3c-3cosc
=4(cos3a+cos3b+cos3c)-3(cosa+cosb+cosc)
if x+y+z=0 then x3+y3+z3=3abc
=12cosa cosb cosc